For your attendance check-in to start this week, please write back with your answers to these questions 1) What percent of data in a bell-shaped (normal) distribution is within 1 standard deviation of the mean? 2) What percent of data in a distribution, using Chebyshevs Rule is within 3 standard deviations of the mean? Thanks, Mr. Richardson 8 9 8 6
Answer #1

Sol:ution1:

For bell shaped or normal distributions:

According to empirical rule

68% of the data lies within one standard deviations of the mean.

ANSWER:68%

Solution2:

here k=3

1-1/k^2=1-1/3^2=1-1/9

=0.8888889

=0.8888889*100

=88.89%

>89% of the distribution lies within 3 sd of the mean

ANSWER:89%